The number of monthly breakdown of a computer is a random variable with an average breakdown of 1.8per month. Find the probability that a computer will function for a month:(i). Without breakdown, (ii). With at least one breakdown
Answer by rothauserc(4718)

This problem is solved using the Poisson distribution
:
The mean(u) for the distribution is 1.8
:
P(X) = e^(-u) * u^x/x!
:
(i) x = 0, then
:
P(X=0) = e^(-1.8) * 1.8^0 / 0! = 0.1653
:
(ii) P(at least one breakdown) = 1 - P(X=0) = 1 - 0.1653 = 0.8347
:
Note e = 2.71828 (but use your calculator's e button)
: